HUBBLE SPACE TELESCOPE DAILY REPORT #2842 PERIOD COVERED: 0000Z (UTC) 04/02/01 - 0000Z (UTC) 04/03/01 Daily Status Report as of 093/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 8829 (Observatory Focus Monitor) The WF/PC-2 was used to analyze the HST focus which drifts slowly and shows evidence of undergoing slips of a few microns at random times. The rate of the WFPC2 monitoring program is insufficient to track and/or understand OTA behavior in order to request timely and appropriate secondary mirror corrections. This 14-orbit program obtains a large amount of high signal to noise focus data, sometimes in two Science Instruments at once, and will be sufficient to more accurately define the HST focus. There were no reported problems. 1.2 Completed Three Sets of STIS/CCD 9248 (Probing the Large Scale Structure: Cosmic Shear Observations) The Space Telescope Imaging Spectrograph (CCD) was used to measure the distortion of light bundles from distant galaxies that probe the statistical properties of the intervening inhomogeneous {dark} matter distribution. Its tidal gravitational field distorts the observable image shapes thereby causing a coherent ellipticity pattern {Cosmic Shear}. The observations completed nominally. 1.3 Completed Two Sets of STIS/CCD 8864 (CCD Dark Monitor-Part 2) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the darks for the CCD. The proposal completed nominally. 1.4 Completed Five Sets of WF/PC-2 8816 (Cycle 9 UV Earthflats) The WF/PC-2 was used to obtain sequences of Earth streak flats to improve the quality of pipeline flat fields for the WFPC2 UV filter set and in order to monitor flat field stability. There were no reported problems. 1.5 Completed Three Sets of WF/PC-2 8827 (Cycle 9 Supplemental Darks pt2/3) The WF/PC-2 was used to obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The proposal completed nominally. 1.6 Completed FGS/1 8774 (Calibrating the Mass-Luminosity Relation at the End of the Main Sequence) FGS-1R was used to calibrate the mass-luminosity relation {MLR} for stars less massive than 0.2 Msun, with special emphasis on objects near the stellar/brown dwarf border. Our goals are to determine Mv values to 0.10 magnitude, masses to 5%, and more than double the number of objects with masses determined to be less than 0.20 Msun. The proposal completed with no reported problems. 1.7 Completed Four Sets of WF/PC-2 9244 (POMS Test Proposal: WFII Parallel Archive Proposal Continuation) The WF/PC-2 was used to perform a generic target version of the Archival Pure Parallel program. The program will be used to take parallel images of random areas of the sky, following the recommendations of the Parallels Working Group. The observations completed with no anomalous activity. 1.8 Completed STIS/MA2 8843 (Cycle 9 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA2) was used to perform the routine monitoring of the MAMA detector dark noise. The proposal completed nominally. 1.9 Completed STIS/CCD/MA1 8662 (A Snapshot Survey of the Hot Interstellar Medium) The Space Telescope Imaging Spectrograph (CCD And MA1) was used to obtain snapshot STIS echelle observations of key tracers of hot interstellar gas {C IV, N IV, and Si IV} for selected FUSE Team O VI survey targets with known UV fluxes. By taking advantage of the snapshot observing mode we will efficiently obtain a large number of spectra suitable for the study of the highly ionized hot component of the interstellar medium {ISM}. Our goals are to explore the physical conditions in and distribution of such gas, as well as to explore the nature of the interfaces between the hot ISM and the other interstellar gas phases. The observations completed with no reported problems. 1.10 Completed WF/PC-2/STIS/CCD 8472 (Key Project: Demographics Of Nuclear Black Holes Observations) The WF/PC-2 and the Space Telescope Imaging Spectrograph (CCD) were used to investigate the detection of supermassive black holes {BHs} in normal galaxies which is one of the grand challenges that HST was designed to meet. During this program, the kinematics of nuclear gas disks in the target galaxies will be measured. The proposal completed with no reported problems. 1.11 Completed WF/PC-2 8812 (Cycle 9 Internal Monitor) The WF/PC-2 was used to obtain routine internal monitors for WFPC2 in order to monitor the health of the cameras. A variety of internal exposures are obtained in order to provide a monitor of the integrity of the CCD camera electronics in both bays {gain 7 and gain 15}, a test for quantum efficiency in the CCDs, and a monitor for possible buildup of contaminants on the CCD windows. There were no reported problems. 1.12 Completed Four Sets of WF/PC-2 8811 (Cycle 9 Standard Darks) The WF/PC-2 was used to obtain dark frames every week in order to provide data for the ongoing calibration of the CCD dark current rate, and to monitor and characterize the evolution of hot pixels. Over an extended period these data will also provide a monitor of radiation damage to the CCDs. The proposal completed with no reported problems. 1.13 Completed STIS/CCD 8865 (Bias Monitor-Part 2) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the bias in the 1x1, 1x2, 2x1, and 2x2 bin settings at gain=1, and 1x1 at gain = 4, to build up high-S/N superbiases and track the evolution of hot columns. The proposal completed with no anomalous activity. 1.14 Completed STIS/CCD 8591 (The Smallest Nuclear Black Holes) The Space Telescope Imaging Spectrograph (CCD) was used to observe small nuclear black holes which are the last major unexplored part of BH parameter space, searching for the smallest BHs that HST can possibly find. The proposal completed with no reported anomalies. 1.15 Completed Four Sets of STIS/CCD 8851 (Sparse Field CTE test {Cycle 9}) The Space Telescope Imaging Spectrograph (CCD) was used to perform CTE measurements that are made using the "sparse field test", along both the serial and parallel axes. The proposal completed with no reported anomalies. 1.16 Completed Two Sets of WF/PC-2 8577 (Ozone, Condensates, and Dust in the Martian Atmosphere) The WF/PC-2 was used to study the spatial and seasonal variations in ozone, condensates, and dust in the Martian atmosphere. The data obtained promise to be critical in addressing recent breakthroughs in understanding the basic radiative, transport, and microphysical processes that provide for both long-term and short-term balance within the global Mars climate system. There were no reported problems. 1.17 Completed WF/PC-2 8773 (Expansion Parallax Distances to Planetary Nebulae) The WF/PC-2 was used to observe planetary nebulae {PNe} for which the lack of reliable distances is especially troubling. Acquired by generalized statistical methods in all but a handful of cases, individual PNe distances are often uncertain by factors of two or more. The observations completed nominally. 1.18 Completed WF/PC-2 8583 (Imaging Snapshots of Asteroids) The WF/PC-2 was used to obtain images of the fifty largest main belt asteroids that have favorable apparitions during cycle 9. The images will be searched for companion bodies, as well as mineralogical variegation on the resolved main bodies. There were no reported problems. 1.19 Completed WF/PC-2 8599 (A Census of Nuclear Star Clusters in Late-Type Spiral Galaxies) The WF/PC-2 was used to conduct an I-band snapshot survey of a well-defined sample of nearby, face-on spiral galaxies of type Scd or later. The proposal completed nominally. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 8 Successful: 8 Scheduled Re-acquisitions: 8 Successful: 8 2.2 FHST Updates: Scheduled: 20 Successful: 19 The update at 092/1146Z failed. 2.3 Operations Notes: Using ROP SR-1A, the SSR EDAC error counter was cleared three times. Engineering status buffer parameters were adjusted at 092/1202Z per ROP DF-18A. 486 table dumps were performed between 092/1542Z and 092/1635Z per ROP DF-1A. A TTR was written for a two-minute loss of engineering data beginning at 092/2035Z due to an incorrect MA frequency scheduled pass. Another TTR was generated for a required re-transmit at 093/032934Z during a NSSC-1 load. The ephemeris tables were uplinked at 093/0431Z as directed by ROP DF-07A. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /CAW